Please solve
Solve each of the following equations, correct to two decimal places:
(a) 6𝑒−2𝑥 = 0.62 (b) 5ln 4𝑥 = 9.84

Answer by Boreal(15235)

6e-0.62=2x=15.69
x=7.84
-
5 ln 4x=9.84
ln 4x=1.968
e of both sides
4x=7.156
x=1.79
check
ln 7.16=1.968
5 ln 7.16=9.84